# ZeroCAP: Zero-Shot Multi-Robot Context Aware Pattern Formation via Large Language Models

### v2

@@ I Introduction
Pattern formation, a key idea in the organization of multi-robot systems (MRS) [1], involves arranging robots in a specific way to reach a common objective. This ability is critical for applications such as environmental surveillance, precision farming, cooperative transportation, and emergency management. Its value lies in coordinating robots to perform tasks beyond the capabilities of individuals, enhancing effectiveness, adaptability, and robustness. In this paper, we go beyond traditional geometric arrangements, focusing on purposeful pattern formations-where patterns are closely tied to user provided contexts and environments. This approach marks an evolution in MRS, steering the field towards more interactive and object-centric tasks that can be applicable to real world scenarios as shown in Fig 1.
@@ V-A Baseline Comparisons
The system is benchmarked against three baselines. The first is an RL-based approach with hand-crafted rewards, focusing on pattern formation without natural language comprehension. The second and third baselines utilize GPT-4 Vision [34] and LLava-v1.6 [35], relying solely on vision-language models to complete tasks. The LLMs used in our experiments include GPT-4 [36], llama70b [37], and Claude [38], allowing for performance comparison across models and baselines. In ablation studies, GPT-4 is the default LLM configuration for ZeroCAP.
@@ VI Results and Analysis
Task-wise, general pattern formation proved the most challenging, likely due to the abstract nature of the instructions, requiring nuanced interpretation. Infill tasks, with clearer objectives, achieved the highest SR and GCR, particularly with GPT-4, as they mainly involved placing robots within a specified region. Caging tasks, although more structured, still posed difficulties, but ZeroCAP managed consistent barrier formations effectively.
@@ VII Conclusion
This work presents ZeroCAP, a framework integrating LLMs with MRS for zero-shot, context-aware pattern formation. By translating natural language instructions into precise spatial arrangements, ZeroCAP overcomes VLM spatial localization limitations by decoupling spatial reasoning from visual processing, using edge and vertex representations for more accurate pattern formation.
Currently, due to VLMs limitations in processing 3D data, being trained mostly on 2D images [11], this work is constrained to a 2D framework to ensure robustness. As VLMs evolve, future exploration could extend to 3D. The system is deployable with a map or overhead camera and could be expanded using multiple cameras in the future for wider scale and coverage. While our focus has been on static formations, dynamic formations that adapt to real-time changes hold significant promise. Further investigation of complex environments with partial occlusions could also improve robustness. This early-stage research lays the groundwork for more intuitive robotic coordination with implications in applications such as surveillance and logistics.

### v3

@@ I Introduction
Pattern formation, a key idea in the organization of multi-robot systems (MRS) [1], involves arranging robots in a specific way to reach a common objective. This ability is critical for applications such as environmental surveillance, precision farming, cooperative transportation, and emergency management. Its value lies in coordinating robots to perform tasks beyond the capabilities of individuals, enhancing effectiveness, adaptability, and robustness. In this paper, we go beyond traditional geometric arrangements, focusing on purposeful pattern formations-where patterns are closely tied to user provided contexts and environments. This approach marks an evolution in MRS, steering the field towards more interactive and object-centric tasks that can be applicable to real world scenarios as shown in Fig. 1.
@@ V-A Baseline Comparisons
The system is benchmarked against three baselines: an RL-based approach with hand-crafted rewards for pattern formation—where Caging rewarded enclosing a target, Infill incentivized structured spacing inside a region, and General used a proximity-based reward for clustering robots near a reference point—and two vision-language models, GPT-4 Vision [34] and LLava-v1.6 [35]. Experiments employ GPT-4 [36], LLaMA-70B [37], and Claude [38], with GPT-4 serving as the default in ablation studies. Unlike traditional rule-based pattern formation methods, ZeroCAP interprets natural language instructions to enable context-aware formations, rendering direct comparisons with classical approaches infeasible due to their lack of user intent integration.
@@ VI Results and Analysis
Task-wise, general pattern formation was the most challenging due to language ambiguity and the lack of structured references, leading to misinterpretations in robot placements. Infill tasks achieved the highest SR and GCR, though occasional failures occurred from uneven spacing inside the region. Caging tasks, while more structured, encountered difficulties with precise localization, often resulting in incomplete barriers due to gaps in robot placement, but ZeroCAP managed consistent barrier formations effectively.
@@ VII Conclusion
This work presents ZeroCAP, a framework integrating LLMs with MRS for zero-shot, context-aware pattern formation. By translating natural language instructions into precise spatial arrangements, ZeroCAP overcomes VLM spatial localization limitations by decoupling spatial reasoning from visual processing, using edge and vertex representations for more accurate pattern formation. Currently, as VLMs are mostly trained on 2D images [11], ZeroCAP is confined to a 2D framework for robustness, though future developments may extend it to 3D. The system is deployable with a map or overhead camera and could be expanded using multiple cameras in the future for wider coverage. While our focus has been on static formations, dynamic, real-time adaptations in complex, occluded environments hold significant promise, paving the way for more intuitive robotic coordination in applications such as surveillance and logistics.